OpenStudy (mendicant_bias):

I'm trying to figure out how to evaluate a triple integral, posted below in a second. The first integration step is easy, but the second, I'm not so sure about.

OpenStudy (mendicant_bias):

\[\int\limits_{0}^{2}\int\limits_{0}^{4-x^{2}}\int\limits_{0}^{x} \frac {\sin(2z)}{4-z}dydzdx\]

OpenStudy (mendicant_bias):

Integrating first with respect to y, you get ysin(2z)/(4-z); integrating w.r.t. z is where I get caught here. Any ideas/hints?

OpenStudy (mendicant_bias):

Whoops, and evaluating y, it just turns into x. Should've gone further. But that still doesn't help me on the second and third integral.

ganeshie8 (ganeshie8):

\[\int\limits_{0}^{2}\int\limits_{0}^{4-x^{2}} \frac {x\sin(2z)}{4-z} dzdx\]

ganeshie8 (ganeshie8):

change it to dxdz

OpenStudy (mendicant_bias):

Oh, okay, derp. Didn't even think about changing order of integration. One moment.

OpenStudy (mendicant_bias):

\[\int\limits_{0}^{2}\int\limits_{0}^{4x^{2}} \frac {xsin(2z)}{4 - z} = \int\limits_{0}^{2}\int\limits_{0}^{\sqrt{z/4}}\]

OpenStudy (mendicant_bias):

(Same integrand other than switching the variables of itnegration, but I'm worried about that top left, outermost integral's upper bound. I'm not sure how to find it, it's obviously a constant and I can intuitively *think* it's two, but I'm not so sure about proving that. Anyways, I'll run with that for now unless corrected and integrate.)

OpenStudy (mendicant_bias):

\[\frac {x^{2}\sin(2z)}{2(4 - z)} = \frac {zsin(2z)}{8(4-z)}\]

ganeshie8 (ganeshie8):

you need to sketch the bounds first

ganeshie8 (ganeshie8):

sketch the region bounded by below curves : z = 4-x^2 z = 0 x = 0 x = 2

OpenStudy (mendicant_bias):